Form a partial differential equation by eliminating the functions f and g from z =

yf(x) + xg(y).


1
Expert's answer
2022-03-03T04:35:12-0500

Given, "z = yf(x) + xg(y)."


Differentiating partially with respect to "x" and "y" respectively, we get


"\begin{aligned}
p &= \dfrac{\partial z}{\partial x} = yf'(x)+g(y) \qquad (1)\\
q &= \dfrac{\partial z}{\partial y} = f(x)+xg'(y) \qquad (2)\\ 
\end{aligned}"


Differentiating equation (1) with respect to y, we get

"s= \dfrac{\partial^{2} z}{\partial y \partial x} = f'(x) + g'(y) \qquad(3)"


Multiplying equation (1) by x, Multiplying equation (2) by y, and adding we get


"\begin{aligned}
xp +yq &= xyf'(x)+xg(y) + yf(x)+xyg'(y) \\
&= xy(f'(x)+g'(y)) + yf(x)+xg(y)\\
&= xys+z \qquad(\text{Using equation (3) and}~ z = yf(x) + xg(y))\\
\therefore z &=xp + yq - xys 
\end{aligned}"

which is the required partial differential equation.